OpenStudy (anonymous):

what is the value of y in this equation y-14=6

OpenStudy (anonymous):

y=6+14

OpenStudy (anonymous):

which is 20

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

y-14=6 y=14+6 y=20